# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

### v1

@@ 7. Power integral bases and canonical number systems in number fields
Number systems and their generalizations have been intensively studied for a long time. Here we present an important generalization for the number field case, point out its close connection with power integral bases and formulate an application of the above Corollary 5.3 to this generalization.
@@
$\mathcal{N}(\alpha)=\{0,1,\ldots,|N_{K/\mathbb{Q}}(\alpha)-1\}$
@@
$a_{0}+a_{1}\alpha+\cdots+a_{k}\alpha^{k}\text{ with }a_{i}\in\mathcal{N}(\alpha)\text{ for }i=0,\ldots,k,\,a_{k}\neq 0$
@@
Kovács (1981) proved the following fundamental result.
Theorem 7.1 (Kovács, 1981).
@@
Together with the above Corollary 5.3 this implies that it is effectively decidable whether there exists a CNS in $\mathcal{O}_{K}$. Corollary 5.3 provides even a general algorithm to determine all power integral bases in $\mathcal{O}_{K}$. Using this, Kovács and Pethő (1991) characterized the bases of all CNS’s of $\mathcal{O}_{K}$. As a consequence they showed as follows.
Theorem 7.2 (Kovács and Pethő, 1991).
@@
In fact, they extended their result to any order of $K$ as well.
We note that Brunotte (2001) considerably improved the procedure of Kovács and Pethő (1991) and gave an efficient algorithm for finding all such CNS’s, provided that one has an efficient algorithm for determining all power integral bases in $\mathcal{O}_{K}$. As was seen in Section 6, such an efficient algorithm is known for number fields $K$ of degree at most $6$ if their discriminants are not too large in absolute value.
For surveys on CNS’s, we refer to Brunotte (2001), Pethő (2004), Brunotte, Huszti and Pethő (2006), Evertse, Győry, Pethő and Thuswaldner (2019) and the references given there.
8. Generalizations and further applications
For a detailed, comprehensive treatment of the reduction theory of integral polynomials with given non-zero discriminant, including Theorems 4.1, 4.2, 4.4 and their generalizations and applications, we refer to the monograph Evertse and Győry (2017) and the references given there. For the algorithmic aspect one can consult also the books Smart (1998) and Gaál (2019), and for general results over integral domains of characteristic 00 that are finitely generated over $\mathbb{Z}$, the book Evertse and Győry (2022).
Below we present, resp. mention some of the most important generalizations and applications.
8.1. Generalizations
@@
Theorems 4.1 and 4.4 were proved in Birch and Merriman (1972) resp. in Evertse and Győry (1991) in more general forms, over number fields, more precisely over the $S$-integers of a number field. Theorem 4.2, (ii) was extended to this more general situation in Győry (1978a, 1981).
We present a generalization of Corollary 5.3 in the relative case. Let $L$ be a number field, $K/L$ a finite relative extension, $\mathcal{O}_{L}$ and $\mathcal{O}_{K}$ the rings of integers of $L$ and $K$. Then $K/L$ is called monogenic if $\mathcal{O}_{K}=\mathcal{O}_{L}[\alpha]$ for some $\alpha\in\mathcal{O}_{K}$, and $\{1,\alpha,\ldots,\alpha^{n-1}\}$ is a relative power integral basis of $K$ over $L$, where $n=[K:L]$, i.e., an $\mathcal{O}_{L}$-module basis of $\mathcal{O}_{K}$. We say that $\alpha,\beta\in\mathcal{O}_{K}$ are $\mathcal{O}_{L}$-equivalent if $\beta=a+\varepsilon\alpha$ for some $a\in\mathcal{O}_{L}$ and unit $\varepsilon$ in $\mathcal{O}_{L}$. If $\alpha$ is a generator of $\mathcal{O}_{K}$ over $\mathcal{O}_{L}$, i.e., $\mathcal{O}_{K}=\mathcal{O}_{L}[\alpha]$ then so is every $\beta$ $\mathcal{O}_{K}$-equivalent to $\alpha$.
Theorem 8.1 (Győry, 1978a).
There are only finitely many $\mathcal{O}_{L}$-equivalence classes of $\alpha\in\mathcal{O}_{K}$ with $\mathcal{O}_{K}=\mathcal{O}_{L}[\alpha]$, and a full set of representatives of such $\alpha$ can be, at least in principle, effectively determined.
For the best known version with explicit bounds, see Corollary 8.4.13 in Evertse and Győry (2017).
We now present two general finiteness theorems where the ground ring is an integrally closed integral domain $A$ of characteristic 00 that is finitely generated over $\mathbb{Z}$ as a $\mathbb{Z}$-algebra, i.e., $A=\mathbb{Z}[z_{1},\ldots,z_{r}]$, where we allow some of the $z_{i}$ to be transcendental. The unit group of $A$ is denoted by $A^{*}$.
We say that the monic polynomials $f,g\in A[X]$ are $A$-equivalent if $g(X)=f(X+a)$ with some $a\in A$. Then $f$ and $g$ have the same discriminant.
Theorem 8.2 (Győry, 1982).
Let $G$ be a finite extension of the quotient field of $A$. Up to $A$-equivalence, there are only finitely many monic $f(X)$ in $A[X]$ with given non-zero discriminant $D$ having all their zeros in $G$.
This was made effective by Győry (1984) in a special case, and in full generality by Evertse and Győry (2017), provided that $A,G$ and $D$ are given effectively in the sense defined in Evertse and Győry (2017, 2022).
Theorem 8.3 (Evertse and Győry, 2017, 2022).
Up to $A$-equivalence, there are only finitely many monic $f$ in $A[X]$ with $D(f)=D$, and if $A,G,D$ are effectively given, all these $f$ can be effectively determined.
Problem 2.
Are these statements true without fixing the splitting field $G$?
  • Theorem 4.4 was later generalized for decomposable forms in more than two variables in Evertse and Győry (1992) and Győry (1994).
  • Theorem 4.2, (ii) and Corollaries 5.4, 5.6 were extended to the case when $D$ resp. $I$ is replaced by $p_{1}^{z_{1}},\ldots,p_{s}^{z_{s}}$, where $p_{1},\ldots,p_{s}$ are fixed primes and $z_{1},\ldots,z_{s}$ unknown non-negative integers; see Győry (1978b, 1981), Trelina (1977a, 1977b), Győry and Papp (1977). These results yielded e.g. explicit lower bound for the greatest prime factor of discriminant and index of an integer of a number field. For generalizations for the number field case, see Győry (1980a).
  • Corollary 5.4 was generalized for more general discriminant form equations
    $D(x_{2}\omega_{2}+\cdots+x_{k}\omega_{k})=D\text{ in }x_{2},\ldots,x_{k}\in\mathbb{Z}$
    where $1,\omega_{2},\ldots,\omega_{k}$ are $\mathbb{Q}$-linearly independent elements of a number field $K$ of degree $n$ with $k\leq n$; see e.g. Győry (1976) and Győry (2000) with a further generalization for the number field case.
    (8.1) $F(x_{1},\ldots,x_{m})=F\text{ in }x_{1},\ldots,x_{m}\in\mathbb{Z},$
  • Let $K$ be a number field with ring of integers $\mathcal{O}_{K}$, and $D\neq 0$ an integer. It is a special case of Corollary 5.1 that, up to $\mathbb{Z}$-equivalence, the equation
    (8.2) $D(\alpha)=D\text{ in }\alpha\in\mathcal{O}_{K}$
    has only finitely many solutions, and that all of these can be effectively determined.
    Let $A=\mathbb{Z}[z_{1},\ldots,z_{r}]$ be an integral domain of characteristic 00 with algebraic or transcendental generators $z_{1},\ldots,z_{r}$, $L$ its quotient field, and $\Omega$ a finite étale $L$-algebra (i.e. a direct product of finite extensions $K_{1},\ldots,K_{t}$ of $L$). Denote by $A_{\Omega}$ the integral closure of $A$ in $\Omega$. The discriminant of $\alpha\in A_{\Omega}$ over $L$ with $\Omega=L[\alpha]$ is given by $D_{L}(\alpha):=D(f_{\alpha})$, where $f_{\alpha}$ is the monic minimal polynomial of $\alpha$ over $L$.
    Let $\mathcal{O}$ be an $A$-order of $\Omega$, i.e. an $A$-subalgebra of $A_{\Omega}$ which spans $\Omega$ as an $L$-vector space.
    Slightly inconsistently with earlier given definitions, we say that $\alpha,\beta\in\mathcal{O}$ are $A$-equivalent if $\beta-\alpha\in A$. One verifies that if $\alpha,\beta\in\mathcal{O}$ are $A$-equivalent then $f_{\alpha}$, $f_{\beta}$ are $A$-equivalent, and thus, $D_{L}(\beta)=D_{L}(\alpha)$.
    Let $D$ be a non-zero element of $L$. Consider the following generalization of equation (8.2):
    (8.3) $D_{L}(\alpha)=D\text{ in }\alpha\in\mathcal{O}.$
    For an integral domain $B$, denote by $B^{+}$ the additive group of $B$.
    Theorem 8.4 (Evertse and Győry, 2022).
    If
    (8.4) $(\mathcal{O}\cap L)^{+}/A^{+}\text{{ is finite,}}$,
    then the set of $\alpha\in\mathcal{O}$ with (8.3) is a union of finitely many $A$-equivalence classes. Moreover, if $A,\Omega,\mathcal{O}$ and $D$ are given effectively in a well-defined way, one can determine a set consisting of precisely one element from each of these classes.
    The condition (8.4) is necessary and decidable.
8.2. Further applications
a
Theorems 4.1, 4.2 and 4.4 as well as their various versions and generalizations led to many applications. Some of them were treated in Sections 4 to 7. Below we briefly present some others in their simplest form. For further applications, we refer to the survey papers Győry (1980a, 2000, 2006), the books Győry (1980b), Smart (1998), Evertse and Győry (2017, 2022), Gaál (2019), and the references given there. Applications to classical Diophantine equations.
@@ 8.2. Further applications
    Using various variants of Theorem 4.2 (ii), Trelina (1985) and, for $n=3,m=2$, Pintér (1995) gave effective upper bounds for $|y|$ that depend on $m,n$ and $|D(f)|$, but not on the height of $f$. We recall that the height of $f$ can be arbitrarily large with respect to $|D(f)|$. Furthermore, Győry and Pintér (2008) showed that for each solution $x,y$ of (8.5) with ${\rm gcd}(y,D(f))=1$, $|y|^{m}$ can be effectively bounded in terms of the radical of $D(f)$, i.e. the product of the distinct prime factors of $D(f)$. Brindza, Evertse and Győry (1991), Haristoy (2003) and Győry and Pintér (2008) gave upper bounds even for $m$ that depend only on $n$ and $|D(f)|$.
  • Let $F\in\mathbb{Z}[X,Y]$ be an irreducible binary form of degree $n\geq 3$ and discriminant $D$, let $p_{1},\ldots,p_{s}\,\,(s\geq 0)$ be distinct primes at most $P$, and let $m$ be a positive integer coprime with $p_{1},\ldots,p_{s}$. There are several upper bounds for the number of solutions $x,y$ of the Thue equation
@@
    where $z_{1},\ldots,z_{2}$ are also unknown non-negative integers.
@@
    Using a quantitative version of Theorem 4.4, i.e. the general effective result of Evertse and Győry (1991a) on binary forms of given degree and given discriminant, previously obtained upper bounds for the number of solutions of these equations were substantially improved under the assumptions that $n,D,m,s$ and $P$ satisfy some additional conditions. Such improved upper bounds were derived in Stewart (1991) for (8.8) with ${\rm gcd}(x,y)=1$ when $m>C_{1}$, in Brindza (1996) for (8.6) with ${\rm gcd}(x,y)=1$ when $m>C_{2}$, and in Thunder (1995) for (8.7) when $m>C_{3}$, where $C_{1},C_{2},C_{3}$ are effectively computable numbers such that $C_{1}$ depends on $n,|D|,P,s$ and $C_{2},C_{3}$ on $n$ and $|D|$. Further, Evertse and Győry (1991b) showed that if $|D|>C_{4}$, then the number of coprime solutions of (8.7) is at most $6n$ if $n>400$, and by Győry (2001) it is at most $628n+$ 6 if $|D|>C_{5}$ and $3\leq n\leq 400$. For $m=1$ and $|D|>C_{6}$, this was later improved by Akhtari (2012) to $211n-$ 2. Here $C_{4},C_{5},C_{6}$ are effectively computable numbers such that $C_{4},C_{5}$ depend on $m$ and $n$, and $C_{6}$ on $n$. Together with a quantitative version of Theorem 4.4, i.e. with the result of Evertse and Győry (1991a), these imply that for given $n\geq 3$ and $m\geq 1$, there are only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes of irreducible binary forms $F\in\mathbb{Z}[X,Y]$ of degree $n$ for which the number of coprime solutions of (8.7) exceeds $628n+$ 6 or $211n-$ 2 if $m=1$.
@@
  • The quantitative version of Theorem 4.4, proved in Evertse and Győry (1991a) was also applied in Evertse (1993) to bound the number of solutions of some resultant inequalities, and in Ribenboim (2006) to binary forms with given discriminant, having additional conditions on the coefficients. We remark that using the improved and completely explicit version Theorem 4.4* of Evertse and Győry (2017), the above quoted applications can be made more precise.
Further applications of Theorems 4.2 and 4.4:
@@
  • Some applications of Theorem 4.2 were given to the reducibility of a general class of polynomials of the form $g(f(X))$ where $f,g$ are monic polynomials, $g(X)$ is irreducible with CM splitting field. For given prime $p$, there are up to $\mathbb{Z}$-equivalence only finitely many $f\in\mathbb{Z}[X]$ of degree $p$ with distinct real zeros for which $g(f(X))$ is reducible; see Győry (1976, 1982).

### v2

@@ 7. Power integral bases and canonical number systems in number fields
Number systems and their generalizations have been intensively studied for a long time. Here we present an important generalization for the number field case, point out its close connection with power integral bases and formulate an application of the above Theorem 5.5 to this generalization.
@@
$\mathcal{N}(\alpha)=\{0,1,\ldots,|N_{K/\mathbb{Q}}(\alpha)|-1\}$
@@
$a_{0}+a_{1}\alpha+\cdots+a_{k}\alpha^{k}\text{ with }a_{i}\in\mathcal{N}(\alpha)\text{ for }i=0,\ldots,k,\,a_{k}\neq 0.$
@@
B. Kovács (1981) proved the following fundamental theorem.
Theorem 7.1 (B. Kovács, 1981).
@@
Together with the above Theorem 5.5 of Győry (1976) this implies that it is effectively decidable whether there exists a CNS in $\mathcal{O}_{K}$. Theorem 5.5 provides even a general algorithm to determine all power integral bases in $\mathcal{O}_{K}$. Using this, B. Kovács and Pethő (1991) proved as follows.
Theorem 7.2 (B. Kovács and Pethő, 1991).
@@
In fact, using Theorem 5.6, they extended their result to any order $\mathcal{O}$ of $K$ as well. In an order $\mathcal{O}$, a canonical number system $\{\alpha,\mathcal{N}(\alpha)\}$ is defined in a similar way as in $\mathcal{O}_{K}$.
We note that Brunotte (2001) considerably improved the procedure of B. Kovács and Pethő (1991) and gave an efficient algorithm for finding all such CNS’s, provided that one has an efficient algorithm for determining all power integral bases in $\mathcal{O}_{K}$, resp. in $\mathcal{O}$. As was seen in Section 6, such an algorithm is known for number fields $K$ of degree at most $4$ if their discriminants are not too large in absolute value.
B. Kovács and Pethő (1991) gave also a complete, effective characterization of CNS’s in number fields and in their orders.
Theorem 7.3 (B. Kovács and Pethő, 1991).
Let $\mathcal{O}$ be an order in a number field $K$. There exist $\alpha_{1},\ldots,\alpha_{t}\in\mathcal{O}$, $n_{1},\ldots,n_{t}\in\mathbb{Z}$, $N_{1},\ldots,N_{t}$ finite subsets of $\mathbb{Z}$, which are all effectively computable, such that $\{\alpha,\mathcal{N}(\alpha)\}$ is a CNS in $\mathcal{O}$, if and only if $\alpha=\alpha_{i}-h$ for some integers $i,h$ with $1\leq i\leq t$ and $h\geq n_{i}$ or $h\in N_{i}$.
Several generalizations and applications have been obtained. Pethő and Varga (2017) generalized the result of B. Kovács to CNS’s over imaginary quadratic Euclidean domains. Pethő and Thuswaldner (2018) study CNS’s in relative extensions. Most of the results of B. Kovács and Pethő (1991) are generalized to this situation. Further generalizations are in Evertse, Győry, Pethő and Thuswaldner (2019) over general orders.
Pethő (1991) introduced the notion of CNS polynomials. The monic polynomial $P(X)\in\mathbb{Z}[X]$ is called CNS polynomial if $|P(0)|\geq 2$ and for every $0\neq Q(X)\in\mathbb{Z}[X]$ there exist unique integers $\ell\geq 0$, $q_{1},\ldots,q_{\ell}\in\{0,1,\ldots,|P(0)|-1\}$ such that
$Q(X)\equiv\sum_{j=0}^{\ell}q_{j}x^{j}\pmod{P(X)}$.
He proved that if $P(X)$ is irreducible and monic and $\alpha$ is one of the zeros of $P(X)$, then $P(X)$ is a CNS polynomial if and only if $\{\alpha,0,1,\ldots,|P(0)|-1\}$ is a CNS in $\mathbb{Z}[\alpha]$.
A. Kovács (2001) computed all CNS polynomials with $P(0)=2$ up to degree $8$. This computation was extended up to degree $14$ in Burcsi and A. Kovács (2008).
Akiyama, Borbély, Brunotte, Pethő and Thuswaldner (2005) defined the shift radix system (SRS). It is a discrete dynamical system, which is a common generalization of CNS polynomials and some kind of $\beta$ representations of real numbers. Many properties of SRS were also described.
For surveys, we refer to Brunotte (2001), Pethő (2004), Brunotte, Huszti and Pethő (2006), Komornik (2011), Evertse, Győry, Pethő and Thuswaldner (2019) and the references given there.
8. Further consequences and applications of the reduction theory
The main results from the effective reduction theory for polynomials discussed before, i.e., Theorems 4.2 and 4.4, as well as their various versions led to many applications. Some of them were treated in Sections 4 to 7. Below we briefly present some others in their simplest form. For further applications, we refer to the survey paper Győry (2006), the books Győry (1980b), Smart (1998), Evertse and Győry (2017) and the references given there.
8.1. Applications to classical Diophantine equations.
@@ 8. Further consequences and applications of the reduction theory
Theorem 4.2 can be applied to superelliptic equations and the Schinzel–Tijdeman equation.
@@
    Applying various variants of Theorem 4.2 to the polynomial $f$ and then using Baker’s method for the reduced equation, Trelina (1985) and, for $n=3,m=2$, Pintér (1995) gave effective upper bounds for $|y|$ that depend on $m,n$ and $|D(f)|$, but not on the height of $f$. We recall that the height of $f$ can be arbitrarily large with respect to $|D(f)|$. Furthermore, Győry and Pintér (2008) showed that for each solution $x,y$ of (8.1) with ${\rm gcd}(y,D(f))=1$, $|y|^{m}$ can be effectively bounded in terms of the radical of $D(f)$, i.e. the product of the distinct prime factors of $D(f)$. It should be noted that $|D(f)|$ can be arbitrarily large with respect to its radical. Brindza, Evertse and Győry (1991), Haristoy (2003) and Győry and Pintér (2008) gave upper bounds even for $m$ that depend only on $n$ and $|D(f)|$.
  • Consider now an application of Theorem 4.2 to equations of discriminant type
    (8.2) $D(x_{1},\ldots,x_{n})=D\text{ in }x_{1},\ldots,x_{n}\in\mathbb{Z},$
    where $D(x_{1},\ldots,x_{n}):=D(f(X))$ is the discriminant of the polynomial $f(X)=X^{n}+x_{1}X^{n-1}+\cdots+x_{n}$ in $X$, and $D\neq 0$ is a given rational integer. If $(x_{1},\ldots,x_{n})$ is a solution of (8.2) then so is
    $(x_{1}^{\ast},\ldots,x_{n}^{\ast})=\left(\frac{f^{(n-1)}(a)}{(n-1)!},\ldots,f(a)\right)\text{ for any }a\in\mathbb{Z},$
    where $X^{n}+x_{1}^{\ast}X^{n-1}+\cdots+x_{n}^{\ast}=:f^{\ast}(X)=f(X+a)$. Such a set of solutions of (8.2) is called a family of solutions. Using a quantitative version of his Theorem 4.2, Győry (1976) proved that (8.2) has only finitely many families of solutions and a representative of every family can be effectively determined. Theorem 4.2* gives a considerable improvement of this result of Győry (1976).
The binary form variant of Theorem 4.4 can be applied to Thue equations, Thue inequalities and Thue–Mahler equations.
  • Let $F\in\mathbb{Z}[X,Y]$ be an irreducible binary form of degree $n\geq 3$ and discriminant $D$, let $p_{1},\ldots,p_{s}\,\,(s\geq 0)$ be distinct primes not exceeding $P$, and let $m$ be a positive integer coprime with $p_{1},\ldots,p_{s}$. There are several upper bounds for the number of solutions $x,y$ of the Thue equation
@@
    where $z_{1},\ldots,z_{s}$ are also unknown non-negative integers.
@@
    Using a quantitative binary form version of Theorem 4.4, e.g. the general effective Theorem 1 of Evertse and Győry (1991a) on binary forms of given degree and given discriminant over $\mathbb{Z}$, the previously obtained upper bounds for the number of solutions of these equations were substantially improved under the assumptions that $n,D,m,s$ and $P$ satisfy some additional conditions. Such improved upper bounds were derived in Stewart (1991) for (8.5) with ${\rm gcd}(x,y)=1$ when $m>C_{1}$, in Brindza (1996) for (8.3) with ${\rm gcd}(x,y)=1$ when $m>C_{2}$, and in Thunder (1995) for (8.4) when $m>C_{3}$, where $C_{1},C_{2},C_{3}$ are effectively computable numbers such that $C_{1}$ depends on $n,|D|,P,s$ and $C_{2},C_{3}$ on $n$ and $|D|$. Further, Evertse and Győry (1991b) showed that if $|D|>C_{4}$, then the number of coprime solutions of (8.4) is at most $6n$ if $n>400$, and by Győry (2001) it is at most $28n+6$ if $|D|>C_{5}$ and $3\leq n\leq 400$. For $m=1$ and $|D|>C_{6}$, this was later improved by Akhtari (2012) to 11n211n-2. Here $C_{4},C_{5},C_{6}$ are effectively computable numbers such that $C_{4},C_{5}$ depend on $m$ and $n$, and $C_{6}$ on $n$. Together with the above mentioned quantitative version of Theorem 4.4, these imply that for given $n\geq 3$ and $m\geq 1$, there are only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes of irreducible binary forms $F\in\mathbb{Z}[X,Y]$ of degree $n$ for which the number of coprime solutions of (8.4) exceeds $28n+6$ or 11n211n-2 if $m=1$.
@@
  • The quantitative version of Theorem 4.4, proved in Evertse and Győry (1991a) was also applied in Evertse (1993) to bound the number of solutions of some resultant inequalities, and in Ribenboim (2006) to binary forms with given discriminant, having additional conditions on the coefficients.
    We remark that using the improved and completely explicit version Theorem 4.4* of Evertse and Győry (2017), the above quoted applications can be made more precise.
8.2. Some other applications of Theorems 4.2 and 4.4
@@
  • Some applications of Theorem 4.2 were given to the reducibility of a general class of polynomials of the form $g(f(X))$ where $f,g$ are monic polynomials, $g(X)$ is irreducible with CM splitting field. For given prime $p$ and $g\in\mathbb{Z}[X]$, there are up to $\mathbb{Z}$-equivalence only finitely many $f\in\mathbb{Z}[X]$ of degree $p$ with distinct real zeros for which $g(f(X))$ is reducible; see Győry (1976, 1982).
  • In Evertse and Győry (1991a), a quantitative binary form variant of Theorem 4.4 was utilized to give effective upper bounds for the minimal non-zero absolute value of binary forms at integral points.